# Phase Transitions in the Fluctuations of Functionals of Random Neural Networks

### 21 Apr 2026 v1

@@ 1.1 Informal statements of the main results
In our context, $Q$ is always $\geq 2$. As anticipated, the answer to 1.1 is fully determined by the fixed-point structure of $\kappa$. A first distinction based on the nature of fixed points was studied in [19], where three regimes were considered: the low-disorder regime $k^{\prime}(1)<1$; the high-disorder regime $k^{\prime}(1)>1$ and the (critical) sparse regime $k^{\prime}(1)=1$. Then, it was shown that $\kappa^{\prime}(1)$ governs an important connection between the geometric properties of the excursion sets of a random neural network and the rate of decay of its angular power spectrum. In this paper, we will further explore the nature of the fixed points of $\kappa$, and we will show how the answer to 1.1 depends on higher order features that go beyond the derivative at $1$.
@@ 1.1.1 Low-disorder: limiting fields and limiting functionals
In the low-disorder regime, we will introduce some regularity conditions on the behaviour of the covariance function at the origin (see 1 below). Under these conditions, we shall prove that $\widehat{\kappa}_{L}\rightarrow\widehat{\kappa}_{\infty}$ uniformly in the compact sets of [1,1][-1,1], where $\hat{\kappa}_{\infty}$ is the covariance function of a well-defined limiting Gaussian field $\widehat{T}_{\infty}$. Moreover, we have (see Theorem 3.3)
@@
The distribution of $F_{\infty}$ is generally not Gaussian, see Remark 3.4.
@@ 1.1.2 Sparse/high-disorder: sharp phase transition to non-Gaussian fluctuations
In the other two regimes, we shall consider different technical assumptions: 1 in the high-disorder case and 2 in the sparse case. Here we prove that $\widehat{\kappa}_{L}$, once suitably rescaled, converges (pointwise) on (1,1)(-1,1) to a limit function $\widehat{\kappa}_{\infty}$, which is not the covariance function of a limit Gaussian field on $\mathbb{S}^{d}$ as in the low-disorder case. Indeed, we have
@@ 1.2.1 Large-domain asymptotics on $\mathbb{Z}^{d}$ and $\mathbb{R}^{d}$
1.2.2 High-frequency asymtptotics
@@ 1.2.3 Our framework: deep asymptotics
Let us fix a centred, stationary Gaussian field $B=(B(x))_{x\in\mathbb{R}^{d}}$ with unit variance and assume (1.4). Observe that, after the change of variable $x=Lx^{\prime}$, addressing 1.1 for functionals defined on the increasing domain $LD$ is equivalent to considering a fixed domain $D\subseteq\mathbb{R}^{d}$ and a sequence of Gaussian fields $(B_{Lx})_{x\in\mathbb{R}^{d}}$ with covariance $\rho(L\cdot)$. The corresponding sequence of functionals to be studied is as follows.
@@
To conclude, we observe another interesting parallelism, related to the short/long-range dependence, see again Subsection 1.2.1. In both cases, taking two disjoint compact sets $D_{1},D_{2}$ (in $\mathbb{R}^{d}$ or $\mathbb{S}^{d}$), one could prove (see, e.g. [39] for large-domain asymptotics and arguing as in Corollary 4.3 for deep asymptotics) that the functionals $F_{L}(D_{1})$, $F_{L}(D_{2})$ are asymptotically uncorrelated (short-range dependence) if $\alpha\geq d/Q$ and asymptotically correlated (long-range dependence) if $\alpha<d/Q$. These analogies suggest the presence of a general structure that may emerge when limiting functions and their singularities are involved. However, the analysis of this general framework is beyond the scope of the present paper.
@@ 3 Main results
Our goal is to answer 1.1, i.e., understand the limit in distribution of $\widetilde{F}_{L}$ as $L\rightarrow\infty$, in the three regimes discovered in [19]:
@@
Under 1 and $\kappa^{\prime}(1)<1$, we find that $\widehat{T}_{L}$ converges to a Gaussian field $\widehat{T}_{\infty}$ in $L^{2}(\Omega\times\mathbb{S}^{d})$ and
@@
The fact that $Z_{Q}(\beta)$ is well defined in $L^{2}(\Omega)$ when (3.7) is satisfied is proved in Lemma 4.8. We also denote $\widetilde{Z}_{Q}(\beta):=Z_{Q}(\beta)\,\operatorname{Var}(Z_{Q}(\beta))^{-1/2}$ if $\operatorname{Var}(Z_{Q}(\beta))\neq 0$.
@@
In this case, one can show that (see (2.15) and Lemma 4.8)
@@
If (3.9) does not hold, the previous representation does not work, but $Z_{Q}(\beta)$ is still in the $Q$-th chaos and equal to $I_{q}\Big(\mathcal{G}_{q}\sqrt{|g_{\beta}|}^{\otimes q}\Big)$ if (3.7) holds, see Lemma 4.8.
We are now ready to state the main result in the high-disorder regime. Recall that here we assume $\varphi$ non-odd when $\kappa$ is odd (see Remark 3.5).
@@
Assume 1 and $\kappa^{\prime}(1)>1$. Then $\kappa$ admits a unique fixed point $b\in(-1,1)$, satisfying $b\geq 0$ and $\kappa^{\prime}(b)\in[0,1)$. Assume that $\kappa^{\prime}(b)\neq 0$ and set
@@
Let us move to the sparse regime. In this case, the iterates of the kernel converge to a non-hyperbolic fixed point. The absence of hyperbolicity prevents a first-order analysis; consequently, stronger regularity assumptions are required at this point, beyond 1.
@@
This assumption is not very restrictive for the activation functions that are commonly adopted. It is satisfied, for instance, by the covariance kernels associated with ReLU and ReLU-like activations (see Remark 4.17); moreover, it covers all “regular” activations (see Remark 4.18).
@@
Under 2, let $\nu=\gamma_{1}-1$. Then, there exist $\beta_{0}$ and $\beta_{1}$ (depending only on $\kappa$) such that
@@
converge to a non-trivial limiting function $\beta(u)$ on (1,1)(-1,1). See Lemma 4.11 and Lemma 4.19.
@@ 4 Proofs
Moreover, if $\kappa\in C^{1}([-1,1])$ and admits the representation (4.1), then it can also be written in the form (2.1) with $\sigma\in\mathbb{D}^{1,2}$. In particular, for a NNGF $T_{L}$ with activation $\sigma$ and kernel $\kappa$ of the form  (2.1), satisfying 1 or 2 as in Theorem 3.3, Theorem 3.8 and Theorem 3.11, we have that $\kappa\in C^{1}$, $\sigma\in\mathbb{D}^{1,2}$ and the above properties are all fulfilled.
@@ 4.1 Low-disorder regime
Under 1 and $\kappa^{\prime}(1)<1$, for every $x\in[-1,1]$, we have
@@
where the limit is strictly positive because $\beta$ is negative in (1,1)(-1,1) by Lemma 4.1. Thus we have
@@ 4.2 Asymptotic analysis in the sparse and high-disorder regimes
High-disorder. One can prove that $\kappa$ admits a unique fixed point $b\in(-1,1)$ and that $b\geq 0$ with $\kappa^{\prime}(b)\in[0,1)$ (see Lemma 4.11 for more details). The case $\kappa^{\prime}(b)=0$, possible only if $b=0$, will be always excluded in our theorems, see Remark 4.13. We use the following notation
@@
Sparse. Let $\gamma_{1}$ as in 2 then we denote with
@@
See Lemmas 4.11 and 4.19 and the corresponding proofs in Sections 5.2 and 5.6 for more details.
Using the unified notation introduced above, we will prove the following result in both the high-disorder regime (see Corollary 4.15) and the sparse regime (see Corollary 4.22)
@@
Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or 2. Then the following asymptotics holds for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
Let $F_{L}[q]$ be the $q$th chaotic component of $F_{L}$ (cf. (3.6)). Under the same assumptions of Proposition 4.2, we have the following asymptotics for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
Now, the claim follows by Proposition 4.2. ∎
@@ 4.3 Central limit theorem in the critical case
Let $F_{L}=F_{L}(\varphi)$ be as in (3.1), with Hermite rank $Q$, $F_{L}[Q]$ as in (3.6) and $\varphi$ non-odd if $\kappa$ is odd. Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or 2. If $Q=d/(2\nu)$, then
@@
We first analyze the variance. From the assumption on the symmetry of $\kappa$ and $\varphi$ and since $Q=d/(2\nu)$, one can use Corollary 4.3 and hence
@@
where the second-last inequality follows by $\sum_{q}q!\varphi_{q}^{2}<\infty$ (since $\varphi\in L^{2}(\mathbb{R},e^{-x^{2}/2})$) and the last ones by  Proposition 4.2. In particular,
@@
Using Proposition 2.1 we have
@@
Now, using Lemma 4.4 we obtain
@@
where the last inequality follows from Proposition 4.2. Using $\operatorname{Var}(F_{L}[Q])\asymp v_{L}^{d/(2\nu)}\log(v_{L}^{-1})$, we obtain
@@ 4.4 Central limit theorem in the non-critical case
Let $F_{L}=F_{L}(\varphi)$ be as in (3.1), with Hermite rank $Q$, $F_{L}[Q]$ as in (3.6) and $\varphi$ non-odd if $\kappa$ is odd. Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or 2. If $Q>d/(2\nu)$, then
@@
Hence, by Proposition 4.2 and  Corollary 4.3 we have $\operatorname{Var}(F_{L})\asymp v_{L}^{d/(2\nu)}$.
@@
where $\widetilde{F}_{L,N_{L}}$ is as in Lemma 4.7 (replacing $N$ with $N_{L}$). To bound the first term in (4.8), one can use Lemma 4.7. Indeed,
@@
where the second-last inequality follows since $N_{L}\geq Q$ and the last follows by Proposition 4.2 (with $c>0$ independent of $Q$ and $L$). Using Proposition 2.1 we have
@@
Now, using Proposition 4.2 and since $|\widehat{\kappa}_{L}|\leq 1$, we obtain, for some $c>0$ independent of $q$ and $L$ (WLOG $r\leq Q-r$)
@@ 4.5 Non-central limit theorem
In this section, we prove the non-central limit theorem in the regime $Q<d/(2\nu)$. We begin by showing that Definition 3.6 is well posed.
@@
Let $F_{L}=F_{L}(\varphi)$ be as in (3.1), with Hermite rank $Q$, $F_{L}[Q]$ as in (3.6) and $\varphi$ non-odd if $\kappa$ is odd. Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or 2. If $Q<d/(2\nu)$, then
@@
Let us start from the variance. From the assumption on the symmetry of $\kappa$ and $\varphi$ and since $Q<d/(2\nu)$, one can use Corollary 4.3 and hence
@@
By Hermite isometry, Lemma 4.8 and recalling the puntual convergence $\frac{g_{L}}{v_{L}}\to g_{\infty}$ (cf. (4.7)), we can show that
@@
if we prove that $R_{L}\to R_{\infty}<\infty$, the claim follows by  (4.7) and generalized dominated convergence. Since $\widehat{\beta}$ is the pointwise limit of $\widehat{\kappa}_{L}$, by Proposition 5.10 below also $\widehat{\beta}$ satisfies the hypothesis of Lemma 4.9. Thus, since
@@
using Lemma 4.9 we obtain
@@
We conclude using Proposition 4.2. ∎
In the following two subsections, we prove  Proposition 4.2 and  (4.7) in both cases, high-disorder and sparse.
4.6 Proof of Proposition 4.2 and (4.7) in the high-disorder regime
@@
Let $\kappa$ be such that $\kappa^{\prime}(1)>1$. Assume that 1 holds. Then there exists a unique fixed point $b\in(-1,1)$, satisfying $b\geq 0$ and $\kappa^{\prime}(b)\in[0,1)$. Moreover, $\kappa^{\prime}(b)=0$ is possible only if $b=0$. Assume that $\lambda:=\kappa^{\prime}(b)\neq 0$ and set
@@
Let $\beta_{L}$, $\beta$, $\lambda$ be as defined in Lemma 4.11. Then, we have the following asymptotics for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
Case $\bm{q\geq\frac{d}{2\nu}}$. Let $r=1-c_{b}$, where $c_{b}$ is defined in Lemma 4.11. The asymptotic behaviour of the integral is entirely driven by the singularity of $\beta$ at $u=1$. Accordingly, we write
@@
Using the previous proposition, we have the following corollary, that is  Proposition 4.2 in the high-disorder case.
@@
Let $\widehat{\kappa}_{L}$ be as in (3.3), $\widehat{\beta}$ as in Proposition 4.2 and assume $\kappa^{\prime}(1)>1$, $\lambda=\kappa^{\prime}(b)\neq 0$ and 1. Then, we have the following asymptotics for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
If $q<d/(2\nu)$, then $\beta\in L^{k}(\mu)$ also for $k<q$ and so the claim follows using Proposition 4.12 and (4.23). Let us now prove that $\int\widehat{\beta}^{q}\,\mathrm{d}\mu=0$ $\iff$ $\kappa$ and $q$ are both odd. Since $\widehat{\beta}\in L^{2}(\mu)$ and we have the $L^{2}(\mu)$-decomposition
@@
Moreover, by Lemma 4.11 and (4.24), the sequence $\beta_{L}-\int\beta_{L}\,\,\mathrm{d}\mu$ converges pointwise to $\widehat{\beta}$ on (1,1)(-1,1). By (P0)(P3), each $\beta_{L}$ is real analytic on (1,1)(-1,1) and satisfies $\beta_{L}^{(k)}(0)\geq 0$ for all $k\in\mathbb{N}$. Hence, by Proposition 5.10, the limit $\widehat{\beta}$ is also real analytic on (1,1)(-1,1) and satisfies $\widehat{\beta}^{(k)}(0)\geq 0$ for all $k\in\mathbb{N}$. As a consequence, applying Lemma 4.9 and recalling (4.7), we obtain
@@
Conversely, assume that $\kappa$ is not odd. Then $\widehat{\beta}$ is not odd. [^7] If $\widehat{\beta}$ were odd, then $\widehat{\beta}(-u)=-\widehat{\beta}(u)$ would contradict the boundary behaviour implied by Lemma 4.11. Using the parity property $G_{\ell,d}(-u)=(-1)^{\ell}G_{\ell,d}(u)$ together with (4.26), it follows that there exists at least one even index $\ell$ such that $g_{\infty}(\ell)>0$. Therefore, by (2.7) and the lower bound (4.27), we conclude that $\int\widehat{\beta}^{q}\,\,\mathrm{d}\mu>0$.
If $q\geq d/(2\nu)$, then by Proposition 4.12, for every $1\leq k\leq q-1$ (WLOG $k\leq q-k$), we get
@@
The proof of (4.7) when $\beta\in L^{2}(\mu)$, i.e. when $2<d/2\nu$, mirrors the proof of the convergence of $\frac{g_{L}(\ell,m)}{\kappa^{\prime}(1)^{L}}$ in the low-disorder case, but using Lemma 4.11 and dominated convergence.
4.7 Proof of Proposition 4.2 and (4.7) in the sparse regime
The proof of these cases is more similar to the proofs in the high disorder, but with different velocity and a different control for the integrals. First, we observe that 2 is satisfied for activation functions commonly considered in the literature.
@@
Thus, the ReLU kernel satisfies 2 with $\gamma_{2}>2(\gamma_{1}-1)$
@@
Let $\kappa\in C^{4}([-1,1])$ be a normalized kernel associated with some activation functions. Then $\kappa$ satisfies 2 with $\gamma_{2}=2(\gamma_{1}-1)$. Indeed, by Taylor expansion, we have
@@
The next lemma is analogous to Lemma 4.11 and the proof is deferred to Section 5.6

### 23 Apr 2026 v2

@@ 1.1 Informal statements of the main results
In our context, $Q$ is always $\geq 2$. As anticipated, the answer to Question 1.1 is fully determined by the fixed-point structure of $\kappa$. A first distinction based on the nature of fixed points was studied in [19], where three regimes were considered: the low-disorder regime $k^{\prime}(1)<1$; the high-disorder regime $k^{\prime}(1)>1$ and the (critical) sparse regime $k^{\prime}(1)=1$. Then, it was shown that $\kappa^{\prime}(1)$ governs an important connection between the geometric properties of the excursion sets of a random neural network and the rate of decay of its angular power spectrum. In this paper, we will further explore the nature of the fixed points of $\kappa$, and we will show how the answer to Question 1.1 depends on higher order features that go beyond the derivative at $1$.
@@ 1.1.1 Low-disorder: limiting fields and limiting functionals
In the low-disorder regime, we will introduce some regularity conditions on the behaviour of the covariance function at the origin (see Assumption 1 below). Under these conditions, we shall prove that $\widehat{\kappa}_{L}\rightarrow\widehat{\kappa}_{\infty}$ uniformly in the compact sets of [1,1][-1,1], where $\hat{\kappa}_{\infty}$ is the covariance function of a well-defined limiting Gaussian field $\widehat{T}_{\infty}$. Moreover, we have (see Theorem 3.3)
@@
The distribution of $F_{\infty}$ is generally not Gaussian, see Section 3.
@@ 1.1.2 Sparse/high-disorder: sharp phase transition to non-Gaussian fluctuations
In the other two regimes, we shall consider different technical assumptions: Assumption 1 in the high-disorder case and Assumption 2 in the sparse case. Here we prove that $\widehat{\kappa}_{L}$, once suitably rescaled, converges (pointwise) on (1,1)(-1,1) to a limit function $\widehat{\kappa}_{\infty}$, which is not the covariance function of a limit Gaussian field on $\mathbb{S}^{d}$ as in the low-disorder case. Indeed, we have
@@ 1.2.1 Large-domain asymptotics on $\mathbb{Z}^{d}$ and $\mathbb{R}^{d}$
1.2.2 High-frequency asymptotics
@@ 1.2.3 Our framework: deep asymptotics
Let us fix a centred, stationary Gaussian field $B=(B(x))_{x\in\mathbb{R}^{d}}$ with unit variance and assume (1.4). Observe that, after the change of variable $x=Lx^{\prime}$, addressing Question 1.1 for functionals defined on the increasing domain $LD$ is equivalent to considering a fixed domain $D\subseteq\mathbb{R}^{d}$ and a sequence of Gaussian fields $(B_{Lx})_{x\in\mathbb{R}^{d}}$ with covariance $\rho(L\cdot)$. The corresponding sequence of functionals to be studied is as follows.
@@
To conclude, we observe another interesting parallelism, related to the short/long-range dependence, see again Subsection 1.2.1. In both cases, taking two disjoint compact sets $D_{1},D_{2}$ (in $\mathbb{R}^{d}$ or $\mathbb{S}^{d}$), one could prove (see, e.g. [39] for large-domain asymptotics and arguing as in Section 4.2 for deep asymptotics) that the functionals $F_{L}(D_{1})$, $F_{L}(D_{2})$ are asymptotically uncorrelated (short-range dependence) if $\alpha\geq d/Q$ and asymptotically correlated (long-range dependence) if $\alpha<d/Q$. These analogies suggest the presence of a general structure that may emerge when limiting functions and their singularities are involved. However, the analysis of this general framework is beyond the scope of the present paper.
@@ 3 Main results
Our goal is to answer Question 1.1, i.e., understand the limit in distribution of $\widetilde{F}_{L}$ as $L\rightarrow\infty$, in the three regimes discovered in [19]:
@@
Under Assumption 1 and $\kappa^{\prime}(1)<1$, we find that $\widehat{T}_{L}$ converges to a Gaussian field $\widehat{T}_{\infty}$ in $L^{2}(\Omega\times\mathbb{S}^{d})$ and
@@
The fact that $Z_{Q}(\beta)$ is well defined in $L^{2}(\Omega)$ when (3.7) is satisfied is proved in Section 4.5. We also denote $\widetilde{Z}_{Q}(\beta):=Z_{Q}(\beta)\,\operatorname{Var}(Z_{Q}(\beta))^{-1/2}$ if $\operatorname{Var}(Z_{Q}(\beta))\neq 0$.
@@
In this case, one can show that (see (2.15) and Section 4.5)
@@
If (3.9) does not hold, the previous representation does not work, but $Z_{Q}(\beta)$ is still in the $Q$-th chaos and equal to $I_{q}\Big(\mathcal{G}_{q}\sqrt{|g_{\beta}|}^{\otimes q}\Big)$ if (3.7) holds, see Section 4.5.
We are now ready to state the main result in the high-disorder regime. Recall that here we assume $\varphi$ non-odd when $\kappa$ is odd (see Section 3).
@@
Assume Assumption 1 and $\kappa^{\prime}(1)>1$. Then $\kappa$ admits a unique fixed point $b\in(-1,1)$, satisfying $b\geq 0$ and $\kappa^{\prime}(b)\in[0,1)$. Assume that $\kappa^{\prime}(b)\neq 0$ and set
@@
Let us move to the sparse regime. In this case, the iterates of the kernel converge to a non-hyperbolic fixed point. The absence of hyperbolicity prevents a first-order analysis; consequently, stronger regularity assumptions are required at this point, beyond Assumption 1.
@@
This assumption is not very restrictive for the activation functions that are commonly adopted. It is satisfied, for instance, by the covariance kernels associated with ReLU and ReLU-like activations (see Section 4.7); moreover, it covers all “regular” activations (see Section 4.7).
@@
Under Assumption 2, let $\nu=\gamma_{1}-1$. Then, there exist $\beta_{0}$ and $\beta_{1}$ (depending only on $\kappa$) such that
@@
converge to a non-trivial limiting function $\beta(u)$ on (1,1)(-1,1). See Section 4.6 and Section 4.7.
@@ 4 Proofs
Moreover, if $\kappa\in C^{1}([-1,1])$ and admits the representation (4.1), then it can also be written in the form (2.1) with $\sigma\in\mathbb{D}^{1,2}$. In particular, for a NNGF $T_{L}$ with activation $\sigma$ and kernel $\kappa$ of the form  (2.1), satisfying Assumption 1 or Assumption 2 as in Theorem 3.3, Theorem 3.8 and Theorem 3.11, we have that $\kappa\in C^{1}$, $\sigma\in\mathbb{D}^{1,2}$ and the above properties are all fulfilled.
@@ 4.1 Low-disorder regime
Under Assumption 1 and $\kappa^{\prime}(1)<1$, for every $x\in[-1,1]$, we have
@@
where the limit is strictly positive because $\beta$ is negative in (1,1)(-1,1) by Section 4.1. Thus we have
@@ 4.2 Asymptotic analysis in the sparse and high-disorder regimes
High-disorder. One can prove that $\kappa$ admits a unique fixed point $b\in(-1,1)$ and that $b\geq 0$ with $\kappa^{\prime}(b)\in[0,1)$ (see Section 4.6 for more details). The case $\kappa^{\prime}(b)=0$, possible only if $b=0$, will be always excluded in our theorems, see Section 4.6. We use the following notation
@@
Sparse. Let $\gamma_{1}$ as in Assumption 2 then we denote with
@@
See Sections 4.6 and 4.7 and the corresponding proofs in Sections 5.2 and 5.6 for more details.
Using the unified notation introduced above, we will prove the following result in both the high-disorder regime (see Section 4.6) and the sparse regime (see Section 4.7)
@@
Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either Assumption 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or Assumption 2. Then the following asymptotics holds for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
Let $F_{L}[q]$ be the $q$th chaotic component of $F_{L}$ (cf. (3.6)). Under the same assumptions of Section 4.2, we have the following asymptotics for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
Now, the claim follows by Section 4.2. ∎
@@ 4.3 Central limit theorem in the critical case
Let $F_{L}=F_{L}(\varphi)$ be as in (3.1), with Hermite rank $Q$, $F_{L}[Q]$ as in (3.6) and $\varphi$ non-odd if $\kappa$ is odd. Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either Assumption 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or Assumption 2. If $Q=d/(2\nu)$, then
@@
We first analyze the variance. From the assumption on the symmetry of $\kappa$ and $\varphi$ and since $Q=d/(2\nu)$, one can use Section 4.2 and hence
@@
where the second-last inequality follows by $\sum_{q}q!\varphi_{q}^{2}<\infty$ (since $\varphi\in L^{2}(\mathbb{R},e^{-x^{2}/2})$) and the last ones by  Section 4.2. In particular,
@@
Using Section 2.3 we have
@@
Now, using Section 4.3 we obtain
@@
where the last inequality follows from Section 4.2. Using $\operatorname{Var}(F_{L}[Q])\asymp v_{L}^{d/(2\nu)}\log(v_{L}^{-1})$, we obtain
@@ 4.4 Central limit theorem in the non-critical case
Let $F_{L}=F_{L}(\varphi)$ be as in (3.1), with Hermite rank $Q$, $F_{L}[Q]$ as in (3.6) and $\varphi$ non-odd if $\kappa$ is odd. Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either Assumption 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or Assumption 2. If $Q>d/(2\nu)$, then
@@
Hence, by Section 4.2 and  Section 4.2 we have $\operatorname{Var}(F_{L})\asymp v_{L}^{d/(2\nu)}$.
@@
where $\widetilde{F}_{L,N_{L}}$ is as in Section 4.4 (replacing $N$ with $N_{L}$). To bound the first term in (4.8), one can use Section 4.4. Indeed,
@@
where the second-last inequality follows since $N_{L}\geq Q$ and the last follows by Section 4.2 (with $c>0$ independent of $Q$ and $L$). Using Section 2.3 we have
@@
Now, using Section 4.2 and since $|\widehat{\kappa}_{L}|\leq 1$, we obtain, for some $c>0$ independent of $q$ and $L$ (WLOG $r\leq Q-r$)
@@ 4.5 Non-central limit theorem
In this section, we prove the non-central limit theorem in the regime $Q<d/(2\nu)$. We begin by showing that Section 3 is well posed.
@@
Let $F_{L}=F_{L}(\varphi)$ be as in (3.1), with Hermite rank $Q$, $F_{L}[Q]$ as in (3.6) and $\varphi$ non-odd if $\kappa$ is odd. Let $\widehat{\kappa}_{L}$ be as in (3.3) and assume either Assumption 1, $\kappa^{\prime}(1)>1$ and $\kappa^{\prime}(b)\neq 0$, or Assumption 2. If $Q<d/(2\nu)$, then
@@
Let us start from the variance. From the assumption on the symmetry of $\kappa$ and $\varphi$ and since $Q<d/(2\nu)$, one can use Section 4.2 and hence
@@
By Hermite isometry, Section 4.5 and recalling the puntual convergence $\frac{g_{L}}{v_{L}}\to g_{\infty}$ (cf. (4.7)), we can show that
@@
if we prove that $R_{L}\to R_{\infty}<\infty$, the claim follows by  (4.7) and generalized dominated convergence. Since $\widehat{\beta}$ is the pointwise limit of $\widehat{\kappa}_{L}$, by Section 5.7 below also $\widehat{\beta}$ satisfies the hypothesis of Section 4.5. Thus, since
@@
using Section 4.5 we obtain
@@
We conclude using Section 4.2. ∎
In the following two subsections, we prove  Section 4.2 and  (4.7) in both cases, high-disorder and sparse.
4.6 Proof of Section 4.2 and (4.7) in the high-disorder regime
@@
Let $\kappa$ be such that $\kappa^{\prime}(1)>1$. Assume that Assumption 1 holds. Then there exists a unique fixed point $b\in(-1,1)$, satisfying $b\geq 0$ and $\kappa^{\prime}(b)\in[0,1)$. Moreover, $\kappa^{\prime}(b)=0$ is possible only if $b=0$. Assume that $\lambda:=\kappa^{\prime}(b)\neq 0$ and set
@@
Let $\beta_{L}$, $\beta$, $\lambda$ be as defined in Section 4.6. Then, we have the following asymptotics for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
Case $\bm{q\geq\frac{d}{2\nu}}$. Let $r=1-c_{b}$, where $c_{b}$ is defined in Section 4.6. The asymptotic behaviour of the integral is entirely driven by the singularity of $\beta$ at $u=1$. Accordingly, we write
@@
Using the previous proposition, we have the following corollary, that is  Section 4.2 in the high-disorder case.
@@
Let $\widehat{\kappa}_{L}$ be as in (3.3), $\widehat{\beta}$ as in Section 4.2 and assume $\kappa^{\prime}(1)>1$, $\lambda=\kappa^{\prime}(b)\neq 0$ and Assumption 1. Then, we have the following asymptotics for fixed $q\in\mathbb{N}$, $q\geq 2$.
@@
If $q<d/(2\nu)$, then $\beta\in L^{k}(\mu)$ also for $k<q$ and so the claim follows using Section 4.6 and (4.23). Let us now prove that $\int\widehat{\beta}^{q}\,\mathrm{d}\mu=0$ $\iff$ $\kappa$ and $q$ are both odd. Since $\widehat{\beta}\in L^{2}(\mu)$ and we have the $L^{2}(\mu)$-decomposition
@@
Moreover, by Section 4.6 and (4.24), the sequence $\beta_{L}-\int\beta_{L}\,\,\mathrm{d}\mu$ converges pointwise to $\widehat{\beta}$ on (1,1)(-1,1). By (P0)(P3), each $\beta_{L}$ is real analytic on (1,1)(-1,1) and satisfies $\beta_{L}^{(k)}(0)\geq 0$ for all $k\in\mathbb{N}$. Hence, by Section 5.7, the limit $\widehat{\beta}$ is also real analytic on (1,1)(-1,1) and satisfies $\widehat{\beta}^{(k)}(0)\geq 0$ for all $k\in\mathbb{N}$. As a consequence, applying Section 4.5 and recalling (4.7), we obtain
@@
Conversely, assume that $\kappa$ is not odd. Then $\widehat{\beta}$ is not odd. [^7] If $\widehat{\beta}$ were odd, then $\widehat{\beta}(-u)=-\widehat{\beta}(u)$ would contradict the boundary behaviour implied by Section 4.6. Using the parity property $G_{\ell,d}(-u)=(-1)^{\ell}G_{\ell,d}(u)$ together with (4.26), it follows that there exists at least one even index $\ell$ such that $g_{\infty}(\ell)>0$. Therefore, by (2.7) and the lower bound (4.27), we conclude that $\int\widehat{\beta}^{q}\,\,\mathrm{d}\mu>0$.
If $q\geq d/(2\nu)$, then by Section 4.6, for every $1\leq k\leq q-1$ (WLOG $k\leq q-k$), we get
@@
The proof of (4.7) when $\beta\in L^{2}(\mu)$, i.e. when $2<d/2\nu$, mirrors the proof of the convergence of $\frac{g_{L}(\ell,m)}{\kappa^{\prime}(1)^{L}}$ in the low-disorder case, but using Section 4.6 and dominated convergence.
4.7 Proof of Section 4.2 and (4.7) in the sparse regime
The proof of these cases is more similar to the proofs in the high disorder, but with different velocity and a different control for the integrals. First, we observe that Assumption 2 is satisfied for activation functions commonly considered in the literature.
@@
Thus, the ReLU kernel satisfies Assumption 2 with $\gamma_{2}>2(\gamma_{1}-1)$
@@
Let $\kappa\in C^{4}([-1,1])$ be a normalized kernel associated with some activation functions. Then $\kappa$ satisfies Assumption 2 with $\gamma_{2}=2(\gamma_{1}-1)$. Indeed, by Taylor expansion, we have
@@
The next lemma is analogous to Section 4.6 and the proof is deferred to Section 5.6